An object moves in a straight line starting from rest. There are two stage in The Journey.
1. It gain speed uniformly for 2.0second and attains a 8.0ms-¹
2.it continue at this speed for a further 1.5second. Sketch a graph of speed against time
Find(1)the acceleration in stage (a)
(2)the acceleration in stage (b)
(3)the total distance moved during stages (a)and(b)
1
Expert's answer
2020-10-27T11:16:22-0400

1)


"a=\frac{v}{t}=\frac{8}{2}=4\frac{m}{s^2}"

2)


"a=0\frac{m}{s^2}"

3)


"s=0.5at^2+vt'=0.5(4)(2)^2+(8)(1.5)=20\ m"